# Vibravox: A Dataset of French Speech Captured with Body-conduction Audio Sensors

### v3

@@ IV-B3 Effect of EBEN enhancement in quiet conditions
A supplementary analysis of the PER has been performed, in order to quantify the effect of the bandwidth enhancement from EBEN models trained on speech-clean presented in Subsection IV-A for ASR. This comparison was conducted employing solely the headset phonemizer.
Figure 7: PER on speech-clean of the raw and EBEN-enhanced test recordings phonemized by the headset phonemizer
The results presented as a bar graph in Figure 7 demonstrate that all sensors exhibited a strong improvement in PER thanks to the enhancement obtained with EBEN models. This is a promising outcome, particularly in light of the fact that contemporary speech enhancement systems, when used as a pre-processing layer in ASR systems continue to face challenges in enhancing performance due to the production of artifacts [88]. The aforementioned findings thus corroborate the hypothesis that speech intelligibility is enhanced by EBEN models. However, it is essential to note that the direct parallel between the headset phonemizer and the human ear cannot be taken for granted. Subjective testing could be considered to provide an additional layer of validation.
@@ IV-B4 Experiments in noisy conditions
To complete the analysis of this task, a new set of audio phonemizers was trained on the speech-noisy training set, using the raw, unprocessed samples. Unlike for the speech enhancement task where a clean reference is unavailable in speech-noisy, labels (phonemized transcriptions) remain available as reference data in noisy conditions. Several training strategies were explored, including retraining from the VoxPopuli checkpoint and fine-tuning the previously trained speech-clean models with an even smaller learning rate of $1e^{-6}$. The latter approach demonstrated a clear advantage. Additionally, given that the speech-noisy training set is only 1 hour and 34 minutes in length, an aggressive data augmentation strategy was necessary to prevent overfitting.
Figure 8 presents the test results for phonemizers trained exclusively on the speech-clean subset, along with those refined using speech-noisy training data. These results highlight the impact of signal-to-external-noise ratios on the final performance of the speech-noisy phonemizers, as well as the performance gap between speech-clean and speech-noisy models. Specifically, as microphone-captured noise increases, the task becomes more challenging, and the domain gap between quiet and noisy conditions broadens. However, fine-tuning helps reduce this gap. These findings offer a pseudo-ranking of speech intelligibility in a 90 dB noisy environment across different microphones. Notably, the headset microphone remains the top performer due to its strong directivity toward the speaker’s mouth. Since no processing was applied to reconstruct the bandwidth of body-conducted signals, high-frequency phonemes were lost. A significant performance drop is observed for the in-ear microphone in noisy conditions, likely due to inadequate acoustic sealing for certain ear canal anatomies that do not fit the sensor properly. While a ”one-size-fits-all” design is ideal in theory, molded eartips would provide a better fit in practice, although this would require considerable additional manufacturing time for each participant.
@@ IV-C Speaker verification
This final section addresses speaker verification, which ascertains whether two audio files feature the same speaker, providing a boolean answer. Speaker verification constitutes a subset of speaker recognition tasks [89], which also encompass speaker identification and diarization. This task is especially pertinent in radio communication scenarios with multiple speakers, where BCMs can exacerbate the probability of voice confusion. Such confusion can impede fluid conversations, particularly in noisy environments. Furthermore, robust speaker identification systems are essential in contexts where heavy machinery is operated via voice commands, ensuring rigorous safety protocols are upheld. As stated by [90], ”The presence and use of high-frequency information from the speech signal improves the recognition of individual speakers, particularly in noisy environments.”. Therefore, as highlighted by previous research[91, 92, 93, 94, 95], the lack of these frequencies presents a significant challenge to BCMs speaker verification, necessitating further investigation to produce a more robust technology for these sensors. We therefore assess the efficiency of speaker verification tasks using body-conducted sensors from the Vibravox dataset.
@@ IV-C2 Testing procedure
It is crucial to highlight that the speaker verification task is only composed of a testing procedure, where we utilize a pretrained ECAPA2 model, without any further training. Consequently, we solely utilize the test splits. The speaker verification testing set is constructed by forming pairs of speech signals, designated as $A$ and $B$, in alignment with the methodology outlined in [96]. Notably, $A$ and $B$ may correspond to different airborne or body-conduction microphones. For each speaker, we generate $N_{p}$ pairs with their own recordings and an additional randomly selected $N_{p}$ pairs with recordings from different speakers. To increase the difficulty of the task, we maintain the possibility of separating speakers by gender. Here, $N_{p}$ represents the number of possible pairs for the speaker with the fewest recordings, which is set at 2346. We thus have a total of 98532 test pairs given the fact that the speech-clean test set is composed of 21 speakers. We generate pairs only for the speech-clean subset to evaluate the model under quiet conditions and reuse these same pairs for the noisy condition. The latter is handled through a static mixing of speech-clean with speechless-noisy, which was shown in Section IV-A to be a suitable modeling approach. This allows for a direct comparison of results between the two recording conditions. When comparing pairs of signals, we compute the cosine similarity and the Euclidean distance between their normalized embeddings. After inference on the entire set of pairs, we compute the Equal Error Rate (EER) and the normalized minimum of the Detection Cost Function (DCF) [98] from the cosine similarity scores.
@@ IV-C4 Effect of EBEN enhancement in quiet conditions
To complete this study, the Equal Error Rate (EER) was compared between raw and EBEN-enhanced signals. Contrary to the observations in Section IV-B, a deterioration in performance is evidenced in Figure 10. Note that speaker identity preservation was not included among the three losses—adversarial loss ($\mathcal{L}\mathcal{G}^{adv}$), feature loss ($\mathcal{L}\mathcal{G}^{feat}$), and spectral loss ($\mathcal{L}_{\mathcal{G}}^{spec}$) which were employed to adjust the EBEN parameters. One possible explanation for the observed performance drop is that the ECAPA2 model may interpret the EBEN outputs as being out-of-distribution, which could account for the poor performance. This issue could be addressed in future work by incorporating speaker identity preservation losses into the overall objective, as explored in [99].
Figure 10: EER on speech-clean of the raw and EBEN enhanced recordings with the speaker verification model on pairs of the same audio sensor
@@ V Discussion and Future Work
Building the Vibravox dataset has been a protracted undertaking that has necessitated the achievement of several milestones. These include the construction of equipment, the development of necessary software, the recording of participants, the filtering of data, and other related tasks. We are pleased to announce the release of this highly curated dataset, comprising four subsets: speech-clean, speech-noisy, speechless-clean and speechless-noisy. The total duration of the recordings is over 4 hours, with 188 participants. This corpus is being made available online on HuggingFace via Creative Commons Attribution 4.0 International license. Using Vibravox has proven to yield satisfactory results in speech enhancement, speech-to-phonemes and speaker verification, which are three fundamental speech processing tasks. The EBEN models used for speech enhancement demonstrated the capacity to enhance objective metrics such as Noresqua-MOS and STOI, while simultaneously improving phoneme transcription on every sensor. However, it was found to have a detrimental impact on speaker verification performance. Also, the aforementioned three tasks facilitated an in-depth comprehension of the distinctive acoustic subtleties to the various audio sensors and their respective placements. It is our hope that this release will drive progress in body-conducted speech analysis and facilitate the development of robust communication systems for real-world applications.
@@ Acknowledgements
The authors would like to express their gratitude to all the 200 participants in the Vibravox measurement protocol, whose invaluable contributions made this research possible. We would also like to thank Jean-Baptiste Doc for his assistance in developing 3D-printed components to facilitate the positioning of sensors to the participants’ diverse morphologies, and Philippe Chenevez for his expertise in electrical engineering for the pre-amplifier circuits. This work was granted access to the HPC/AI resources of [CINES / IDRIS / TGCC] under the grant 2022-AD011013469 awarded by GENCI and partially funded by the French National Research Agency under the ANR Grant No. ANR-20-THIA-0002.
@@
In the preparation of this work, the authors utilized DeepL Write and ChatGPT to improve the quality of the English language in the manuscript. Following the use of these tools, the authors thoroughly reviewed and edited the content as necessary, and assume full responsibility for the final version of the publication.
@@ Appendix A Dataset statistics
Appendix B Speech Enhancement training curves on speech-clean subset
@@
Appendix C Speech-to-Phoneme training curves on speech-clean subset
@@ Appendix E Power spectral density of microphones on speechless-clean
Appendix F Speaker Verification EER on speech-clean subset

### v4

@@ IV-B3 Effect of EBEN enhancement in quiet conditions
A supplementary analysis of the PER has been performed in order to quantify the effect of the bandwidth enhancement from EBEN models trained on speech-clean presented in Subsection IV-A for ASR. This comparison was conducted employing solely the headset phonemizer. The results presented as a bar graph in Figure 7 demonstrate that all sensors exhibited a strong improvement in PER thanks to the enhancement obtained with EBEN models. This is a promising outcome, since contemporary speech enhancement systems, when used as a pre-processing layer in ASR systems, continue to face challenges in enhancing performance due to the production of artifacts [90]. The aforementioned findings thus corroborate the hypothesis that speech intelligibility is enhanced by EBEN models. However, it is essential to note that the direct parallel between the headset phonemizer and the human ear cannot be taken for granted. Subjective testing could be considered to provide an additional layer of validation.
Figure 7: PER on speech-clean of raw and EBEN-enhanced test recordings phonemized by the headset phonemizer
@@ IV-B4 Experiments in noisy conditions
To complete the analysis of this task, a new set of audio phonemizers was trained on the speech-noisy training set, using the raw, unprocessed samples. Unlike for the speech enhancement task, where a clean reference is unavailable in speech-noisy, labels (phonemized transcriptions) remain available as reference data in noisy conditions. Several training strategies were explored, including retraining from the VoxPopuli checkpoint and fine-tuning the previously trained speech-clean models with an even smaller learning rate of $1e^{-6}$. The latter approach demonstrated a clear advantage. Additionally, given that the speech-noisy training set is only 1 hour and 34 minutes in length, an aggressive data augmentation strategy was necessary to prevent overfitting.
Figure 8 presents the speech-noisy test results for phonemizers trained exclusively on the speech-clean subset, along with those refined using speech-noisy training data. These results highlight the impact of signal-to-external-noise ratios on the final performance of the phonemizers, as well as the performance gap between speech-clean and speech-noisy models. Specifically, as microphone-captured noise increases, the task becomes more challenging, and the domain gap between quiet and noisy conditions broadens. However, fine-tuning helps reduce this gap. These findings offer a pseudo-ranking of speech intelligibility in a 90 dB noisy environment across different microphones. Notably, the headset microphone remains the top performer due to its strong directivity toward the speaker’s mouth. Since no processing was applied to reconstruct the bandwidth of body-conducted signals, high-frequency phonemes were lost. A significant performance drop is observed for the in-ear microphone in noisy conditions, likely due to inadequate acoustic sealing for certain ear canal anatomies that do not fit the sensor properly. While a ”one-size-fits-all” design is ideal in theory, molded eartips would provide a better fit in practice, although this would require considerable additional manufacturing time for each participant. Throat and temple speech-clean phonemizers remain performant in noise, reflecting the high signal-to-external noise ratio in Table II of those sensors.
@@ IV-C Speaker Verification
This final section addresses speaker verification, which ascertains whether two audio files feature the same speaker, providing a boolean answer. Speaker verification constitutes a subset of speaker recognition tasks [91], which also encompass speaker identification and diarization. This task is especially pertinent in radio communication scenarios with multiple speakers, where BCMs can exacerbate the probability of voice confusion. Such confusion can impede fluid conversations, particularly in noisy environments. Furthermore, robust speaker identification systems are essential in contexts where heavy machinery is operated via voice commands, ensuring rigorous safety protocols are upheld. As stated by [92], ”The presence and use of high-frequency information from the speech signal improves the recognition of individual speakers, particularly in noisy environments.”. Therefore, as highlighted by previous research research[93, 94, 95, 96, 97], the lack of these frequencies presents a significant challenge to BCMs speaker verification, necessitating further investigation to produce a more robust technology for these sensors. We therefore assess the efficiency of speaker verification tasks using body-conducted sensors from the Vibravox dataset.
@@ IV-C2 Testing procedure
It is crucial to highlight that the speaker verification task is only composed of a testing procedure, where we utilize a pretrained ECAPA2 model without any further training. Consequently, we solely utilize the test splits. The speaker verification testing set is constructed by forming pairs of speech signals, designated as $A$ and $B$, in alignment with the methodology outlined in [98]. Notably, $A$ and $B$ may correspond to different airborne or body-conduction microphones. For each speaker, we generate $N_{p}$ pairs with their own recordings and an additional randomly selected $N_{p}$ pairs with recordings from different speakers. To increase the difficulty of the task, we maintain the possibility of separating speakers by gender. Here, $N_{p}$ represents the number of possible pairs for the speaker with the fewest recordings, which is set at 2346. We thus have a total of 98532 test pairs, given the fact that the speech-clean test set is composed of 21 speakers. We generate pairs only for the speech-clean subset to evaluate the model under quiet conditions and reuse these same pairs for the noisy condition. The latter is handled through a static mixing of speech-clean with speechless-noisy, which was shown in Section IV-A to be a suitable modeling approach. This allows for a direct comparison of results between the two recording conditions. When comparing pairs of signals, we compute the cosine similarity and the Euclidean distance between their normalized embeddings. After inference on the entire set of pairs, we compute the Equal Error Rate (EER) and the normalized minimum of the Detection Cost Function (DCF) [100] from the cosine similarity scores.
@@ IV-C4 Effect of EBEN enhancement in quiet conditions
To complete this study, the Equal Error Rate (EER) was compared between raw and EBEN-enhanced signals. Contrary to the observations in Section IV-B, a deterioration in performance is evidenced in Figure 10. Note that speaker identity preservation was not included among the three losses—adversarial loss ($\mathcal{L}_{\mathcal{G}}^{adv}$), feature loss ($\mathcal{L}_{\mathcal{G}}^{feat}$), and spectral loss ($\mathcal{L}_{\mathcal{G}}^{spec}$) which were employed to adjust the EBEN parameters. One possible explanation for the observed performance drop is that the ECAPA2 model may interpret the EBEN outputs as being out-of-distribution, which could account for the poor performance. This issue could be addressed in future work by incorporating speaker identity preservation losses into the overall objective, as explored in [101].
Figure 10: EER on speech-clean of the raw and EBEN-enhanced recordings with the speaker verification model on pairs of the same audio sensor
@@ V Discussion and Future Work
Building the Vibravox dataset has been a protracted undertaking that has necessitated the achievement of several milestones. These include the construction of equipment, the development of necessary software, the recording of participants, the filtering of data, and other related tasks. This highly curated dataset, comprising four subsets —speech-clean, speech-noisy, speechless-clean, and speechless-noisy — is now publicly available. The total duration of the 6-channel recordings is over 45 hours per sensor, with 188 participants. This corpus is being made available online on HuggingFace via Creative Commons Attribution 4.0 International license. Using Vibravox has proven to yield satisfactory results in Speech Enhancement, Speech-to-Phonemes and Speaker Verification, which are three fundamental speech processing tasks. The EBEN models used for speech enhancement demonstrated the capacity to enhance objective metrics such as Noresqa-MOS and STOI while simultaneously improving phoneme transcription on every sensor. However, it was found to have a detrimental impact on speaker verification performance. The aforementioned tasks further contributed to a deeper understanding of the distinctive acoustic subtleties of the various audio sensors and their respective placements.
We hope this release will drive progress in body-conducted speech analysis and support the development of robust communication systems for real-world applications.
@@ Acknowledgements
The authors would like to express their gratitude to the 200 participants in the Vibravox measurement protocol, whose invaluable contributions made this research possible. We would also like to thank Jean-Baptiste Doc for his assistance in developing 3D-printed components that facilitated sensor positioning for participants with diverse morphologies and Philippe Chenevez for his expertise in electrical engineering for the design of pre-amplifier circuits. This work was granted access to the HPC/AI resources of [CINES / IDRIS / TGCC] under the grant 2022-AD011013469 awarded by GENCI and partially funded by the French National Research Agency under the ANR Grant No. ANR-20-THIA-0002.
@@
In the preparation of this work, the authors utilized QuillBot’s Grammar Checker and Mistral’s LeChat to improve the quality of the English language in the manuscript. Following the use of these tools, the authors thoroughly reviewed and edited the content as necessary and assume full responsibility for the final version of the publication.
@@ Appendix A Dataset statistics
Appendix B Speech Enhancement training curves on the speech-clean subset
@@
Appendix C Speech-to-Phonemes training curves on the speech-clean subset
@@ Appendix E Power spectral density of microphones on speechless-clean
Appendix F Speaker Verification EER on the speech-clean subset